OpenStudy (anonymous):

|k/3+2| < or equal to 2

OpenStudy (allank):

You'll have two sets of inequalities here: -(k/3+2)<=2 and k/3+2<=2 Know why?

OpenStudy (anonymous):

its some rule I know for sure teacher thought it but some one told me to subtract two from each side if im right

OpenStudy (allank):

First, know how I split the first expression into two?

OpenStudy (anonymous):

yup

OpenStudy (allank):

Good. Now, the target is to have x on one side of the inequality signs. I'll do the first: -(k/3+2)<=2

OpenStudy (allank):

Multiply the -ve into the brackets: -k/3-2<=2

OpenStudy (allank):

Now add 2 to both sides. -k/3<=4

OpenStudy (allank):

Then comes the part where people make mistakes: Multiply both sides by -3. NOTE: When you multiply or divide both sides by a negative, the INEQUALITY CHANGES DIRECTION: k>=-12

OpenStudy (allank):

Understand?

OpenStudy (anonymous):

mhm thanks

OpenStudy (allank):

you're welcome

OpenStudy (anonymous):

i still have one more question if you dont mind